# Cryptocurrency as an Investable Asset Class: Coming of Age

### v3

@@ 3.3 Stylized fact 3: Mind the Jumps—few factors, higher orders: why less is more
Leibniz wrote that natura non facit saltus—nature makes no jumps. In economics and finance, this often means that most changes occur gradually, with only occasional discontinuities. Indeed, jumps are rare in many asset classes, and this key insight leads Black & Scholes (1973) to model prices as Brownian motions. Cryptocurrencies are different: large, sudden price moves are common. Borri & Santucci de Magistris (2022) show that jumps account for a sizable share of Bitcoin’s return variation and that higher-order moments help explain its risk premium. [^12] See early evidence from Scaillet, Treccani & Trevisan (2020) similarly documenting very frequent jumps in Bitcoin.
@@
This table reports the probability of observing a statistically significant jump (P(J)), the number of such events (N(J)), and quantiles of the ratio of jump variation to realized variance (J/RV), conditional on a significant jump, for Bitcoin, Ethereum, and U.S. stocks. Jumps are identified as the difference between the daily realized variance of Barndorff-Nielsen & Shephard (2002) and the bi-power variation of Barndorff-Nielsen & Shephard (2006). The two realized measures are constructed from tick-level data aggregated at the 5-minute frequency. Significant jumps are identified using the BNS test with a 0.1% significance level. For U.S. equities, values are cross-sectional averages across the constituents of the Dow Jones Industrial Average. Data are from January 1, 2015 to March 31, 2025. The pre-2020 sample is the sample ending Jan 1, 2020. The post-2020 sample is the sample since Jan 1, 2020.
@@
This prevalence of jumps, and the corresponding role of higher-order moments, poses challenges for risk management and portfolio choice (e.g., Borri 2019) and has a direct implication for asset pricing: nonlinearities matter. Motivated by this, a growing literature applies machine-learning methods—especially neural networks—to the cross-section of returns (e.g., Gu, Kelly & Xiu 2020). While neural networks have a universal approximation property (Cybenko 1989; Hornik, Stinchcombe & White 1989), their main drawback in finance is that they are “black box” models: they can fit complex patterns but make it difficult to isolate which economic features drive predicted returns and whether those relationships are stable when the environment changes. [^13] For a review of machine learning in asset pricing, see Nagel (2021).
@@
One alternative to neural nets is the Kolmogorov-Arnold (KA) representation theorem (Kolmogorov 1956; Arnold 1957), introduced in finance by Borri et al. (2024). [^14] Borri et al. (2024), arXiv, April 11, 2024. The theorem shows that a broad class of nonlinear functions can be exactly represented as finite sums of univariate functions composed with a link function—a structure that motivates shallow, low-dimensional architectures (see, e.g., Hecht-Nielsen 1987; Girosi & Poggio 1989; Kurkova 1991). In practice, Borri et al. (2024) implement this idea and construct a single nonlinear factor (the KA factor) that captures a sizable share of the cross-sectional variation in expected returns. [^15] Related work includes Schmidt-Hieber (2021) and Horowitz & Mammen (2007). Subsequent to Borri et al. (2024), Liu et al. (2024) popularize KAN in the machine-learning community, while Toscano, Wang & Karniadakis (2025) discuss computational and scaling limitations.
@@
Another alternative to neural networks follows the traditional finance toolbox—approximate and estimate the stochastic discount factor via a Taylor expansion. However, the Taylor approximation to the higher order generates a very large set of factors, including all of the interactions of the higher-order terms in the Taylor expansion. Borri et al. (2025) develop the properties of a forward-selection Fama-MacBeth (FS-FMB) method for selecting the higher-order factors and their interactions that are most important for the cross-section of returns by sequentially adding one factor or interaction to the selected set at a time and, at each step, choosing the term that maximizes the Fama–MacBeth (FMB) second-step cross-sectional R-squared. This approach is a fully “glass box” model in terms of determining the important factors. [^16] Conceptually, these higher-order interaction terms capture conditional moments, as in systemic-risk measures such as CoVaR (Adrian & Brunnermeier, 2016) and marginal expected shortfall (Acharya et al., 2017). See Bell et al. (2024) for an alternative “glass box” machine learning structure.
@@
Table 5, Panel A, reports second-step FMB results for the crypto-CAPM, or C-1, model and for the C-4 model of Liu, Tsyvinski & Wu (2021). The test assets are the set of 50 portfolios formed using the smart beta strategies described in stylized fact 2 (see Tables 2 and 3), and the CARB strategy described in stylized fact 5 (see Table 6, Panel B). As with the classic CAPM, the C-1 model cannot account for the cross-section of average returns. By contrast, the C-4 model explains 47.3% of the cross-sectional variation, and satisfies the restriction of a zero intercept. Table 5, Panel B, reports the result of a specification which augments the C-4 model with the single KA-factor based on the Kolmogorov-Arnold architecture of Borri et al. (2024). One can think of this factor as a stand-in that captures a variety of different neural nets. The cross-sectional R-squared increases by approximately 25 pp, and the intercept is not statistically different from zero at standard levels. Table 5, Panel C, reports the results from the estimation of the FS-FMB (Borri et al. 2025). At each step, we augment the C-4 by one additional higher-order factor ($h_{j}$) which maximizes the adjusted cross-sectional R-squared. Higher-order factors include the squared and cubed of the factors in the C-4 model, as well their pair-wise interactions up to order three. The procedure selects three higher-order factors. The cross-sectional R-squared of the FS-FMB procedure is 70%, a value very similar to that of the model including the KA-factor. In all models, the estimate of the intercept is not statistically different from zero.
@@
This table reports the results from the second step of the FMB method. Panel A refers to baseline models: the crypto-CAPM (C-1) and the C-4 model of Liu, Tsyvinski & Wu (2022). The table reports the adjusted cross-sectional R-squared, the estimate for the intercept ($\alpha$) and associated t-statistic. The t-statistics are based on Newey-West corrected standard errors. Panel B refers to the C-4 model augmented by the single KA-factor, estimated with the KA procedure of Borri et al. (2024). Panel C refers to the FS-FMB procedure. At each step, we augment the C-4 model by one additional higher-order factor ($h_{j}$) which maximizes the adjusted cross-sectional R-squared. The method stops when the improvement in the R-squared in a given step is smaller than 1 pp.
@@ 3.4 Stylized fact 4: In crypto, the (block)chain drives the gain
Network externalities are a fundamental concept in economics because they shape markets and influence competition, regulation and technology adoption. Classic work has analyzed how networks form and evolve (Jackson & Wolinsky, 1996), how pricing and competition operate on networked platforms (Rochet & Tirole, 2003), and how shocks propagate through financial networks (Allen & Gale, 2000). A central question common in this work is information diffusion and learning in a network (Golub & Jackson 2010; Acemoglu et al. 2011).
@@
Empirically, Liu, Tsyvinski & Wu (2021) show that on-chain user adoption—measured as new user growth—has a statistically and economically significant effect on cryptocurrency prices. Importantly, this single on-chain variable accounts for about 8% of the variation in cryptocurrency returns. In time‑series tests for the largest coins, on-chain user adoption alone can account for more than 10% of the variation in their returns.
Liu, Tsyvinski & Wu (2021) also examine how markets absorb this information and what conditions amplify its relevance. Event‑time evidence shows a sharp price reaction when weekly adoption information is realized, with muted pre‑announcement behavior and no post‑announcement drift—in contrast to equities. The result gives evidence that investors fully incorporate the information at the weekly frequency.
Lastly, Liu, Tsyvinski & Wu (2021) show that the crypto value measure, constructed as the price-to-new-address ratio, negatively predicts cryptocurrency returns—a result we discuss in stylized fact 2.
@@
A central concept in finance is the law of one price—identical securities should sell for identical prices. The law of one price, and the related concept of absence of arbitrage, is important for the existence of a discount factor (Ross 1978; Harrison & Kreps 1979). Deviations from the law of one price have been presented as proof of irrational investor behavior (Lamont & Thaler 2003), limits to arbitrage and market segmentation (Shleifer & Vishny 1997), or previously overlooked risks (Cochrane 2002).
@@
Makarov & Schoar (2020) and Borri & Shakhnov (2023) document sizable, persistent deviations in the dollar price of Bitcoin across exchanges and across fiat- and crypto-quoted pairs. We revisit this evidence on updated data for 510 BTC–fiat pairs traded on 165 centralized exchanges in 49 currencies. Bitcoin price discounts are computed each week relative to the dollar price on Kraken (the benchmark exchange). All quotes are converted to U.S. dollars using daily spot FX rates. In the 2014–2025 sample, the 10th and 90th percentiles of Bitcoin price discounts are –1.8% and 3.7%, respectively. These figures remain virtually unchanged in the post-2020 period (–2.1% and 3.5%). Moreover, discounts are highly persistent, with a half-life of roughly 1.1 weeks in both samples. These deviations—classic examples of market inefficiency and violations of the law of one price—are results of market segmentation induced by capital controls and idiosyncratic domestic shocks that prevent arbitrage.
@@
Fragmentation and limits to arbitrage also have implications for market fragility and flash-crash dynamics. Menkveld & Yueshen (2019) study the May 6, 2010 U.S. equity Flash Crash and argue that a disruption in cross-market arbitrage can make fragmented markets fragile and result in price crashes; they also refer to analogous flash-crash episodes in cryptocurrency markets, including in Ethereum (June 21, 2017) and Bitcoin (October 10, 2017). Because our baseline analysis is conducted at the weekly frequency using close-to-close returns, these very short-lived intraday dislocations are less likely to affect our main results, though they remain relevant for studies of market microstructure and tail risk. [^17] Barbon & Ranaldo (2021) compare centralized and decentralized cryptocurrency exchanges along two key dimensions—market liquidity and price efficiency—and find that decentralized exchanges exhibit lower efficiency.
@@
Table 6, Panel A, reports the properties of an “arbitrage” (CARB) strategy that attempts to exploit these price discrepancies. Each week, we sort BTC–fiat pairs by their discount and form five portfolios. The CARB return for a given pair is defined by buying Bitcoin on Kraken in week $t$ and selling it on the target pair in week t+1t{+}1; portfolio returns are the simple average across constituent pairs. Using weekly formation and holding periods substantially alleviates execution and timing concerns. In practice, capital controls, withdrawal limits, transaction costs, settlement latency, and venue-specific frictions prevent the cross-market trades needed to enforce the law of one price. Not surprisingly, the mean excess returns are implausibly high—on the order of 68.5% per week over 2014–2025, rising to 95% per week from 2020 onward—precisely because the required arbitrage trades are blocked or prohibitively costly. Specifically, Borri & Shakhnov (2022) show that this seemingly exceptional performance is concentrated in the extreme portfolio that holds pairs with the highest Bitcoin price relative to the benchmark quote, where segmentation and frictions are most severe. That is, the cryptocurrency market inefficiencies reflect the classic market segmentation and trading frictions as in Shleifer & Vishny (1997).
@@
As Sophocles writes in Antigone, “Nothing vast enters the life of mortals without a curse.” Early cryptocurrency investors often told a different story: that outsized returns came from cross-exchange “arbitrage.” Public interviews with Sam Bankman-Fried—the founder of the now-defunct exchange FTX—describe how Alameda Research, a hedge fund he helped run, allegedly began by focusing on such trades. [^18] See, for example, Bloomberg’s coverage of the so-called Kimchi premium, which denotes the fact that it was more expensive to buy Bitcoin in Korean Won than in US dollars: https://www.bloomberg.com/news/articles/2024-11-15/south-korea-crypto-boom-what-is-the-kimchi-premium, and Lewis (2023)’s book about Sam Bankman-Fried. However, Borri & Shakhnov (2022) show that implementability is limited to a narrow set of highly liquid pairs on reputable cryptocurrency exchanges without binding capital controls.
@@
We revisit the analysis in Borri & Shakhnov (2022) and construct the CARB strategy for 226 Bitcoin pairs, traded on 96 exchanges against 11 fiat currencies (Australian dollar, Canadian dollar, Danish krone, euro, Hong Kong dollar, Israeli shekel, Japanese yen, Polish złoty, Swiss franc, British pound, and U.S. dollar). These are the pairs that are realistically investable. Table 6, Panel B, documents that excess returns remain statistically significant but are orders of magnitude smaller: the long–short weekly return declines from about $0.68$ to roughly $0.038$ in the full sample, and measures around $0.041$ in the post-2020 sample. In other words, where the strategy is actually implementable, the apparent “free lunch” largely vanishes. These returns are not pure arbitrage—they are compensation for risk. Borri & Shakhnov (2022) show that the expensive pairs (those priced above the benchmark) are systematically riskier—they depreciate more in bad times for crypto-investors, when aggregate liquidity and sentiment are low.
@@
The collapse of Alameda Research and the failure of FTX in November 2022—only a few months after the demise of the synthetic stablecoin ecosystem Terra–Luna described in Liu, Makarov & Schoar (2023), is the most visible manifestation of the results described in this fact. More broadly, the evidence from cryptocurrency markets provides a clean illustration of the classic limits-to-arbitrage view: when trading, settlement, and capital-mobility frictions are economically large, violations of the law of one price can persist rather than being immediately arbitraged away.
@@
Notes: This table reports the mean quintile portfolio returns and long-short portfolio return for the “arbitrage” (CARB) investment strategy that sorts Bitcoin trading pairs by their price deviation from a benchmark Bitcoin price on Kraken. Each week, pairs are sorted into five portfolios. The long-short portfolio is long portfolio 5 and short portfolio 1. The mean returns are the time-series averages of weekly equally-weighted portfolio excess returns. Panel A uses the full sample of 510 BTC–fiat pairs traded on 165 centralized exchanges in 49 currencies. Panel B restricts to 226 BTC–fiat pairs on 96 exchanges against 11 fiat currencies, which are the most liquid fiat currencies on the most reputable exchanges in jurisdictions with minimal capital controls as in Borri & Shakhnov (2022). The post-2020 sample is the sample since Jan 1, 2020. $t$-statistics in parentheses are Newey-West adjusted. *, **, and *** denote significance at the 10%, 5% and 1% levels.

### v4

@@ 3.3 Stylized fact 3: Mind the Jumps—few factors, higher orders: why less is more
Leibniz wrote that natura non facit saltus—nature makes no jumps. In economics and finance, this often means that most changes occur gradually, with only occasional discontinuities. Indeed, jumps are rare in many asset classes, and this key insight leads Black and Scholes (1973) to model prices as Brownian motions. Cryptocurrencies are different: large, sudden price moves are common. Borri and Santucci de Magistris (2022) show that jumps account for a sizable share of Bitcoin’s return variation and that higher-order moments help explain its risk premium. [^11] See early evidence from Scaillet, Treccani, and Trevisan (2020) similarly documenting very frequent jumps in Bitcoin.
@@
This table reports the probability of observing a statistically significant jump (P(J)), the number of such events (N(J)), and quantiles of the ratio of jump variation to realized variance (J/RV), conditional on a significant jump, for Bitcoin, Ethereum, and U.S. stocks. Jumps are identified as the difference between the daily realized variance of Barndorff-Nielsen and Shephard (2002) and the bi-power variation of Barndorff-Nielsen and Shephard (2006). The two realized measures are constructed from tick-level data aggregated at the 5-minute frequency. Significant jumps are identified using the BNS test with a 0.1% significance level. For U.S. equities, values are cross-sectional averages across the constituents of the Dow Jones Industrial Average. Data are from January 1, 2015 to March 31, 2025. The pre-2020 sample is the sample ending Jan 1, 2020. The post-2020 sample is the sample since Jan 1, 2020.
@@
This prevalence of jumps, and the corresponding role of higher-order moments, poses challenges for risk management and portfolio choice (e.g., Borri 2019) and has a direct implication for asset pricing: nonlinearities matter. Motivated by this, a growing literature applies machine-learning methods—especially neural networks—to the cross-section of returns (e.g., Gu, Kelly, and Xiu 2020). While neural networks have a universal approximation property (Cybenko 1989; Hornik, Stinchcombe, and White 1989), their main drawback in finance is that they are “black box” models: they can fit complex patterns but make it difficult to isolate which economic features drive predicted returns and whether those relationships are stable when the environment changes. [^12] For a review of machine learning in asset pricing, see Nagel (2021).
@@
One alternative to neural nets is the Kolmogorov-Arnold (KA) representation theorem (Kolmogorov 1956; Arnold 1957), introduced in finance by Borri et al. (2024). [^13] Borri et al. (2024), arXiv, April 11, 2024. The theorem shows that a broad class of nonlinear functions can be exactly represented as finite sums of univariate functions composed with a link function—a structure that motivates shallow, low-dimensional architectures (see, e.g., Hecht-Nielsen 1987; Girosi and Poggio 1989; Kurkova 1991). In practice, Borri et al. (2024) implement this idea and construct a single nonlinear factor (the KA factor) that captures a sizable share of the cross-sectional variation in expected returns. [^14] Related work includes Schmidt-Hieber (2021) and Horowitz and Mammen (2007). Subsequent to Borri et al. (2024), Liu et al. (2024) popularize KAN in the machine-learning community, while Toscano, Wang, and Karniadakis (2025) discuss computational and scaling limitations.
@@
Another alternative to neural networks follows the traditional finance toolbox—approximate and estimate the stochastic discount factor via a Taylor expansion. However, the Taylor approximation to the higher order generates a very large set of factors, including all of the interactions of the higher-order terms in the Taylor expansion. Borri et al. (2025) develop the properties of a forward-selection Fama-MacBeth (FS-FMB) method for selecting the higher-order factors and their interactions that are most important for the cross-section of returns by sequentially adding one factor or interaction to the selected set at a time and, at each step, choosing the term that maximizes the Fama–MacBeth (FMB) second-step cross-sectional R-squared. This approach is a fully “glass box” model in terms of determining the important factors. [^15] Conceptually, these higher-order interaction terms capture conditional moments, as in systemic-risk measures such as CoVaR (Adrian and Brunnermeier, 2016) and marginal expected shortfall (Acharya et al., 2017). See Bell et al. (2024) for an alternative “glass box” machine learning structure.
@@
Table 5, Panel A, reports second-step FMB results for the crypto-CAPM, or C-1, model and for the C-4 model of Liu, Tsyvinski, and Wu (2021). The test assets are the set of 50 portfolios formed using the smart beta strategies described in stylized fact 2 (see Tables 2 and 3), and the CARB strategy described in stylized fact 5 (see Table 6, Panel B). As with the classic CAPM, the C-1 model cannot account for the cross-section of average returns. By contrast, the C-4 model explains 47.3% of the cross-sectional variation, and satisfies the restriction of a zero intercept. Table 5, Panel B, reports the result of a specification which augments the C-4 model with the single KA-factor based on the Kolmogorov-Arnold architecture of Borri et al. (2024). One can think of this factor as a stand-in that captures a variety of different neural nets. The cross-sectional R-squared increases by approximately 25 pp, and the intercept is not statistically different from zero at standard levels. Table 5, Panel C, reports the results from the estimation of the FS-FMB (Borri et al. 2025). At each step, we augment the C-4 by one additional higher-order factor ($h_{j}$) which maximizes the adjusted cross-sectional R-squared. Higher-order factors include the squared and cubed of the factors in the C-4 model, as well their pair-wise interactions up to order three. The procedure selects three higher-order factors. The cross-sectional R-squared of the FS-FMB procedure is 70%, a value very similar to that of the model including the KA-factor. In all models, the estimate of the intercept is not statistically different from zero.
@@
This table reports the results from the second step of the FMB method. Panel A refers to baseline models: the crypto-CAPM (C-1) and the C-4 model of Liu, Tsyvinski, and Wu (2022). The table reports the adjusted cross-sectional R-squared, the estimate for the intercept ($\alpha$) and associated t-statistic. The t-statistics are based on Newey-West corrected standard errors. Panel B refers to the C-4 model augmented by the single KA-factor, estimated with the KA procedure of Borri et al. (2024). Panel C refers to the FS-FMB procedure. At each step, we augment the C-4 model by one additional higher-order factor ($h_{j}$) which maximizes the adjusted cross-sectional R-squared. The method stops when the improvement in the R-squared in a given step is smaller than 1 pp.
@@ 3.4 Stylized fact 4: In crypto, the (block)chain drives the gain
Network externalities are a fundamental concept in economics because they shape markets and influence competition, regulation and technology adoption. Classic work has analyzed how networks form and evolve (Jackson and Wolinsky, 1996), how pricing and competition operate on networked platforms (Rochet and Tirole, 2003), and how shocks propagate through financial networks (Allen and Gale, 2000). A central question common in this work is information diffusion and learning in a network (Golub and Jackson 2010; Acemoglu et al. 2011).
@@
Empirically, Liu, Tsyvinski, and Wu (2021) show that on-chain user adoption—measured as new user growth—has a statistically and economically significant effect on cryptocurrency prices. Importantly, this single on-chain variable accounts for about 8% of the variation in cryptocurrency returns. In time‑series tests for the largest coins, on-chain user adoption alone can account for more than 10% of the variation in their returns.
Liu, Tsyvinski, and Wu (2021) also examine how markets absorb this information and what conditions amplify its relevance. Event‑time evidence shows a sharp price reaction when weekly adoption information is realized, with muted pre‑announcement behavior and no post‑announcement drift—in contrast to equities. The result gives evidence that investors fully incorporate the information at the weekly frequency.
Lastly, Liu, Tsyvinski, and Wu (2021) show that the crypto value measure, constructed as the price-to-new-address ratio, negatively predicts cryptocurrency returns—a result we discuss in stylized fact 2.
@@
A central concept in finance is the law of one price—identical securities should sell for identical prices. The law of one price, and the related concept of absence of arbitrage, is important for the existence of a discount factor (Ross 1978; Harrison and Kreps 1979). Deviations from the law of one price have been presented as proof of irrational investor behavior (Lamont and Thaler 2003), limits to arbitrage and market segmentation (Shleifer and Vishny 1997), or previously overlooked risks (Cochrane 2002).
@@
Makarov and Schoar (2020) and Borri and Shakhnov (2023) document sizable, persistent deviations in the dollar price of Bitcoin across exchanges and across fiat- and crypto-quoted pairs. We revisit this evidence on updated data for 510 BTC–fiat pairs traded on 165 centralized exchanges in 49 currencies. Bitcoin price discounts are computed each week relative to the dollar price on Kraken (the benchmark exchange). All quotes are converted to U.S. dollars using daily spot FX rates. In the 2014–2025 sample, the 10th and 90th percentiles of Bitcoin price discounts are –1.8% and 3.7%, respectively. These figures remain virtually unchanged in the post-2020 period (–2.1% and 3.5%). Moreover, discounts are highly persistent, with a half-life of roughly 1.1 weeks in both samples. These deviations—classic examples of market inefficiency and violations of the law of one price—are results of market segmentation induced by capital controls and idiosyncratic domestic shocks that prevent arbitrage.
@@
Fragmentation and limits to arbitrage also have implications for market fragility and flash-crash dynamics. Menkveld and Yueshen (2019) study the May 6, 2010 U.S. equity Flash Crash and argue that a disruption in cross-market arbitrage can make fragmented markets fragile and result in price crashes; they also refer to analogous flash-crash episodes in cryptocurrency markets, including in Ethereum (June 21, 2017) and Bitcoin (October 10, 2017). Because our baseline analysis is conducted at the weekly frequency using close-to-close returns, these very short-lived intraday dislocations are less likely to affect our main results, though they remain relevant for studies of market microstructure and tail risk. [^16] Barbon and Ranaldo (2021) compare centralized and decentralized cryptocurrency exchanges along two key dimensions—market liquidity and price efficiency—and find that decentralized exchanges exhibit lower efficiency.
@@
Table 6, Panel A, reports the properties of an “arbitrage” (CARB) strategy that attempts to exploit these price discrepancies. Each week, we sort BTC–fiat pairs by their discount and form five portfolios. The CARB return for a given pair is defined by buying Bitcoin on Kraken in week $t$ and selling it on the target pair in week t+1t{+}1; portfolio returns are the simple average across constituent pairs. Using weekly formation and holding periods substantially alleviates execution and timing concerns. In practice, capital controls, withdrawal limits, transaction costs, settlement latency, and venue-specific frictions prevent the cross-market trades needed to enforce the law of one price. Not surprisingly, the mean excess returns are implausibly high—on the order of 68.5% per week over 2014–2025, rising to 95% per week from 2020 onward—precisely because the required arbitrage trades are blocked or prohibitively costly. Specifically, Borri and Shakhnov (2022) show that this seemingly exceptional performance is concentrated in the extreme portfolio that holds pairs with the highest Bitcoin price relative to the benchmark quote, where segmentation and frictions are most severe. That is, the cryptocurrency market inefficiencies reflect the classic market segmentation and trading frictions as in Shleifer and Vishny (1997).
@@
As Sophocles writes in Antigone, “Nothing vast enters the life of mortals without a curse.” Early cryptocurrency investors often told a different story: that outsized returns came from cross-exchange “arbitrage.” Public interviews with Sam Bankman-Fried—the founder of the now-defunct exchange FTX—describe how Alameda Research, a hedge fund he helped run, allegedly began by focusing on such trades. [^17] See, for example, Bloomberg’s coverage of the so-called Kimchi premium, which denotes the fact that it was more expensive to buy Bitcoin in Korean Won than in US dollars: https://www.bloomberg.com/news/articles/2024-11-15/south-korea-crypto-boom-what-is-the-kimchi-premium, and Lewis (2023)’s book about Sam Bankman-Fried. However, Borri and Shakhnov (2022) show that implementability is limited to a narrow set of highly liquid pairs on reputable cryptocurrency exchanges without binding capital controls.
@@
We revisit the analysis in Borri and Shakhnov (2022) and construct the CARB strategy for 226 Bitcoin pairs, traded on 96 exchanges against 11 fiat currencies (Australian dollar, Canadian dollar, Danish krone, euro, Hong Kong dollar, Israeli shekel, Japanese yen, Polish złoty, Swiss franc, British pound, and U.S. dollar). These are the pairs that are realistically investable. Table 6, Panel B, documents that excess returns remain statistically significant but are orders of magnitude smaller: the long–short weekly return declines from about $0.68$ to roughly $0.038$ in the full sample, and measures around $0.041$ in the post-2020 sample. In other words, where the strategy is actually implementable, the apparent “free lunch” largely vanishes. These returns are not pure arbitrage—they are compensation for risk. Borri and Shakhnov (2022) show that the expensive pairs (those priced above the benchmark) are systematically riskier—they depreciate more in bad times for crypto-investors, when aggregate liquidity and sentiment are low.
@@
The collapse of Alameda Research and the failure of FTX in November 2022—only a few months after the demise of the synthetic stablecoin ecosystem Terra–Luna described in Liu, Makarov, and Schoar (2023), is the most visible manifestation of the results described in this fact. More broadly, the evidence from cryptocurrency markets provides a clean illustration of the classic limits-to-arbitrage view: when trading, settlement, and capital-mobility frictions are economically large, violations of the law of one price can persist rather than being immediately arbitraged away.
@@
Notes: This table reports the mean quintile portfolio returns and long-short portfolio return for the “arbitrage” (CARB) investment strategy that sorts Bitcoin trading pairs by their price deviation from a benchmark Bitcoin price on Kraken. Each week, pairs are sorted into five portfolios. The long-short portfolio is long portfolio 5 and short portfolio 1. The mean returns are the time-series averages of weekly equally-weighted portfolio excess returns. Panel A uses the full sample of 510 BTC–fiat pairs traded on 165 centralized exchanges in 49 currencies. Panel B restricts to 226 BTC–fiat pairs on 96 exchanges against 11 fiat currencies, which are the most liquid fiat currencies on the most reputable exchanges in jurisdictions with minimal capital controls as in Borri and Shakhnov (2022). The post-2020 sample is the sample since Jan 1, 2020. $t$-statistics in parentheses are Newey-West adjusted. *, **, and *** denote significance at the 10%, 5% and 1% levels.